# Entanglement, defects, and $T\bar{T}$ on a black hole background

### v1

@@ 4.2 Entanglement Entropy
In this section we compute the first order correction to the entanglement entropy due to $T\bar{T}$ deformation. We once again consider a radiation subsystem $A\equiv[L,u_{h}]\cup[u_{h},R]$ on a constant time slice in the eternal black hole model discussed previously, with the endpoints of the subsystem $A$ now at $P\equiv(t,L,\rho_{c})$ and $Q\equiv(-t+i2\pi u_{h},R,\rho_{c})$. Here $\rho_{c}$ is a very large yet finite, and denotes the new location of the dual field theory as a result of $T\bar{T}$ deformation. For simplicity of computations, we once again begin in the Poincaré coordinates, where the corresponding endpoints of $A$ are now at $P\equiv(\tau_{1},x_{1},z_{c_{1}})$ and $Q\equiv(\tau_{2},x_{2},z_{c_{2}})$. [^5] Note that is contrast to subsection 3.2, $z_{c_{1}}$ and $z_{c_{2}}$ are now small but finite, and represent the locations of the of the dual field theory due to $T\bar{T}$ deformation. We later transform the final results using eqs. 3.2, 3.4 and 3.6 and express the entanglement entropy in terms of the AdS black string coordinates. Similar to subsection 3.2, we come across two distinct entanglement entropy phases depending on the distance of the subsystem $A$ from the EOW brane.
@@ 4.2.1 Island Phase
Field theory computation : As discussed in the previous section, this phase corresponds to the scenario where the points $P$ and $Q$ are located close to the brane such that the radiation subsystem $A$ has a corresponding island region on the brane. The entanglement entropy in this case may be computed using eq. 2.8, where the effective term may be evaluated using the standard form of the 2-point correlator of twist fields at points $P,P^{\prime},Q$ and $Q^{\prime}$ as
@@
[…]
where the conformal factors $\Omega_{P^{\prime}}$ and $\Omega_{Q^{\prime}}$ are now given by eq. 4.4, and have once again assumed $P^{\prime}\equiv(-y_{{}_{P^{\prime}}},\tau_{{}_{P^{\prime}}})$ and $Q^{\prime}\equiv(-y_{{}_{Q^{\prime}}},\tau_{{}_{Q^{\prime}}})$. Due to the presence of the island region on the brane, the contributions due to the area term in eq. 2.8 must also be taken into consideration. This term may be directly obtained from [76] as
[…]
@@
[…]
where in the final step we utilize the transformations in eqs. 3.2, 3.4 and 3.6 to express the results in the AdS black string coordinates. Note that in the limit $\rho_{c}\to\rho_{\epsilon}$, the dominant term (the first term in subsection 4.2.1) is in exact agreement with the results of the undeformed scenario given by subsection 3.2.1.
Bulk computation : From the bulk perspective the entanglement entropy may be determined using the DES formula in eq. 2.9, where the first term may be evaluated in terms of the geodesic lengths of the two RT surfaces $PP^{\prime}$ and $QQ^{\prime}$ using eq. 3.13 as
[…]
Extremizing the above expression we obtain, upto first order in $z_{c_{1}}$ and $z_{c_{2}}$, $\tau_{{}_{P^{\prime}}}=\tau_{1},y_{{}_{P^{\prime}}}=x_{1}-2z_{c_{1}}e^{\rho_{b}}$ and $\tau_{{}_{Q^{\prime}}}=\tau_{2},y_{{}_{Q^{\prime}}}=x_{2}-2z_{c_{2}}e^{\rho_{b}}$. The final holographic entanglement entropy may then be obtained as subsection 4.2.1 on application of the transformations in eqs. 3.2, 3.4 and 3.6 and the Brown-Henneaux relation.
@@ 4.2.2 No-Island Phase
Field theory computation : This phase corresponds to the scenario where the points $P$ and $Q$ are located far away from the EOW brane, as a result of which the radiation subsystem $A$ has no corresponding entanglement entropy island on the brane. Consequently, the area term in eq. 2.8 does not contribute and the entanglement entropy is equivalent to the effective term in this case, which may be determined in terms of the 2-point correlator of twist fields at $P$ and $Q$ as given in eq. 3.17. The rest of the computation follows as elaborated in subsection 3.2.2, and the final entanglement entropy may then be expressed as
$S(A)=\frac{c}{6}\log\left[\frac{u_{h}}{LR}\left(\Delta_{L}+\Delta_{R}+2\sqrt{\Delta_{L}\Delta_{R}}\cosh\frac{t}{u_{h}}\right)\right]+\frac{c}{3}\rho_{c}.$ (4.12)
@@
Bulk computation : Utilizing the DES formula in eq. 2.9, the holographic entanglement entropy in the no-island phase is simply given by the first term. As elaborated earlier, due to the absence of entanglement entropy islands on the EOW brane the defect term does not contribute in this scenario. The holographic entanglement entropy, which in this case is simply $S_{RT}$, is thus obtained as
$S(A)=S_{RT}=\frac{1}{4G}\cosh^{-1}\left[\frac{(x_{2}-x_{1})^{2}+(\tau_{2}-\tau_{1})^{2}+z_{c_{1}}^{2}+z_{c_{2}}^{2}}{2z_{c_{1}}z_{c_{2}}}\right],$ (4.13)
which matches with eq. 4.12 on application of the transformations described in eqs. 3.2, 3.4 and 4.2 and the Brown-Henneaux relation.
@@ 5 Page curve
In this section we plot and analyse the Page curves for the entanglement entropy of the radiation subsystem $A$, for both undeformed and $T\bar{T}$ deformed scenarios. To this end, we evaluate the entanglement entropy under the assumption $c=6,u_{h}=1,\epsilon_{y}=0.01,L=R=0.5$, and $\rho_{\epsilon}=\rho_{c}=4$. [^6] We assume that $\rho_{c}$ and $\rho_{\epsilon}$ to have same values in this scenario. This is because the objective of this analysis is to investigate the changes in the original Page curve when corrections due to $T\bar{T}$ deformation is taken into account. For this purpose we require the undeformed entanglement entropy to have the same value numerically for both scenarios. In addition we systematically vary $\rho_{b}$ to observe how the Page curve changes with the brane angle. This is illustrated in fig. 6, where the Page curves corresponding to distinct brane angles are depicted using different colors. The Page curve for the deformed scenarios are depicted by the solid curves, while those for the undeformed scenarios are depicted by the dashed curves.
@@
We observe that for both scenarios the entanglement entropy increases with time in the no-island phase. Upon transition to the island phase the entanglement entropy stabilizes to a constant value, which varies with change in the brane angle.
A general trend emerges from fig. 6. We observe in the no-island phase the Page curves overlap for both undeformed and $T\bar{T}$ deformed scenarios and for all values of $\rho_{b}$, indicating that the evolution of the entanglement entropy is independent of the deformation parameter or the brane angle in this phase. However, we observe some interesting features in the island phase. The entanglement entropy in this phase is dependent on both the brane angle and the deformation parameter, with the deformed entanglement entropy being smaller than the undeformed one for a given brane angle. Moreover, the Page time $T_{p}$ (the time of transition between the no-island and the island phases of the entanglement entropy) appears to be dependent on the brane angle as well as the deformation parameter. This dependence can be mathematically explored by deriving the expressions of $T_{p}$ for both the undeformed and the $T\bar{T}$ deformed scenarios as
$T_{p}=\begin{cases}u_{h}\cosh^{-1}\left[\frac{LR\left(e^{2\rho_{b}}+1\right)^{2}-u_{h}\epsilon_{y}^{2}(\Delta_{L}+\Delta_{R})}{2u_{h}\epsilon_{y}^{2}\sqrt{\Delta_{L}\Delta_{R}}}\right]\qquad&\text{Undeformed scenario,}\\ u_{h}\cosh^{-1}\left[\frac{LR\left(e^{2\rho_{b}}+1\right)^{2}e^{-4e^{\rho_{b}-\rho_{c}}}-u_{h}\epsilon_{y}^{2}(\Delta_{L}+\Delta_{R})}{2u_{h}\epsilon_{y}^{2}\sqrt{\Delta_{L}\Delta_{R}}}\right]\qquad&T\bar{T}\text{-deformed scenario,}\end{cases}$ (5.1)
@@
It is interesting to further investigate the dependence of the Page time $T_{p}$ on the brane angle and the deformation parameter and explore any new physics that might emerge. We explore the dependence of $T_{p}$ on the brane angle in fig. 7 for both the undeformed and the $T\bar{T}$ deformed scenarios. Note that though the Page time initially increases with $\rho_{b}$, it rapidly drops to zero for the $T\bar{T}$ deformed scenarios as $\rho_{b}$ further increases. This is in contrast with the undeformed scenario, where the Page time increases consistently with $\rho_{b}$. This behaviour is also observed in [27], where the authors speculate that this rapid decrease in the Page time might be due the CFT being no longer located at the asymptotic boundary at $\rho\to\infty$, but pushed inwards to a finite value in the presence of $T\bar{T}$ deformation.
@@
We further extend the analysis by examining how the Page time changes with the deformation parameter in fig. 8, where we plot $T_{p}$ against $e^{-\rho_{c}}$ (related to the deformation parameter as given in eq. 4.6) for different brane angles. $e^{-\rho_{c}}\to 0$ corresponds to the undeformed scenario, while increasing values indicate $T\bar{T}$ deformed scenarios. We observe that for all brane angles the Page time decreases approximately linearly with the UV cut-off $e^{-\rho_{c}}$, but shows a sudden drop at the end. With increase in $\rho_{b}$, $T_{p}$ starts at a higher value, but decreases much rapidly for high brane angles compared to low brane angles.
@@ 6 Summary and Discussions
In this article, we investigate the proposed duality between the island and the defect extremal surface (DES) prescription for the fine-grained entanglement entropy in a Karch-Randall (KR) braneworld model, where the radiation baths are located in a gravitational background. In particular, we consider an AdS3 black string geometry truncated by an end-of-the-world (EOW) brane for which the effective lower-dimensional perspective consists of a radiation bath defined on an AdS2 black hole background. We consider subsystems on radiation baths corresponding to AdS22 eternal black holes in the lower-dimensional effective picture, and compute the entanglement entropy for distinct phases using the island formula in the effective $2d$ and the DES formula in the bulk picture. We demonstrate an agreement between the two prescriptions, thus establishing that the DES formula is a holographic dual to the island formula. Additionally, the results agree with those existing in the literature obtained using the bulk and the boundary prescription [20, 79], which acts as a check for the accuracy of our computations.
We further investigate the entanglement structure for AdS3 black string geometry with a finite cut-off arising from $T\bar{T}$ deformation [70], truncated by an EOW brane. The usual AdS/CFT duality applies to the Dirichlet boundary on the cut-off surface, while Neumann boundary conditions apply on the EOW brane [76]. We consider an eternal AdS2 black hole coupled to radiation baths (now located on a finite cut-off surface induced by $T\bar{T}$ deformation), and determine the entanglement entropy up to first order for radiation subsystems using both the island and the DES formulas. We observe an agreement between the two prescriptions for both the island and no-island phases of the entanglement entropy, providing another strong consistency check for the proposed duality.
@@
Though [51, 54, 57] have established that modifying the correlator by appropriate insertions of stress tensor components correctly captures the effects of the $T\bar{T}$ deformation, in this article we adopt a different method for our analysis. Similar to [76], we assume that this deformation manifests through modifications in the induced metric defined on the radial cut-off surface and observe an agreement between the results obtained using the island and the DES prescriptions. However, a perturbative theory of $T\bar{T}$ deformed BCFT must still be developed.
This work identifies several promising directions to explore further. An important consideration is analysing the mixed state entanglement structure for both undeformed and deformed scenarios for KR braneworld models. Extending the framework to higher dimensions is an equally crucial advancement, particularly in light of [23], which shows that the DES and the island prescriptions do not align consistently for higher dimensions even for undeformed scenarios. Furthermore, testing the proposed duality between these prescriptions for other deformed theories is also a compelling direction. We leave these open questions for future study.

### v2

@@ 4.2 Entanglement Entropy
In this section we compute the first order correction to the entanglement entropy due to $T\bar{T}$ deformation. We once again consider a radiation subsystem $A\equiv[L,u_{h}]\cup[u_{h},R]$ on a constant time slice in the eternal black hole model discussed previously, with the endpoints of the subsystem $A$ now at $P\equiv(t,L,\rho_{c})$ and $Q\equiv(-t+i2\pi u_{h},R,\rho_{c})$. Here $\rho_{c}$ is a very large yet finite, and denotes the new location of the dual field theory as a result of $T\bar{T}$ deformation. For simplicity of computations, we once again begin in the Poincaré coordinates, where the corresponding endpoints of $A$ are now at $P\equiv(\tau_{1},x_{1},z_{c_{1}})$ and $Q\equiv(\tau_{2},x_{2},z_{c_{2}})$. [^5] Note that is contrast to section 3.2, $z_{c_{1}}$ and $z_{c_{2}}$ are now small but finite, and represent the locations of the of the dual field theory due to $T\bar{T}$ deformation. We later transform the final results using eqs.˜21, 23 and 25 and express the entanglement entropy in terms of the AdS black string coordinates. Similar to section˜3.2, we come across two distinct entanglement entropy phases depending on the distance of the subsystem $A$ from the EOW brane.
@@ 4.2.1 Island Phase
Field theory computation : As discussed in the previous section, this phase corresponds to the scenario where the points $P$ and $Q$ are located close to the brane such that the radiation subsystem $A$ has a corresponding island region on the brane. This is illustrated in fig.˜4. The entanglement entropy in this case may be computed using eq.˜8, where the effective term may be evaluated using the standard form of the 2-point correlator of twist fields at points $P,P^{\prime},Q$ and $Q^{\prime}$ as
@@
$=\frac{c}{6}\left(\log\left[\frac{(y_{{}_{P^{\prime}}}+x_{1})^{2}+(\tau_{1}-\tau_{{}_{P^{\prime}}})^{2}}{z_{c_{1}}\epsilon_{y}}\right]+\log\left[\frac{\ell}{y_{{}_{P^{\prime}}}\operatorname{sech}\rho_{b}+z_{c_{1}}}\right]\right)$
$\qquad+\frac{c}{6}\left(\log\left[\frac{(y_{{}_{Q^{\prime}}}+x_{1})^{2}+(\tau_{2}-\tau_{{}_{Q^{\prime}}})^{2}}{z_{c_{2}}\epsilon_{y}}\right]+\log\left[\frac{\ell}{y_{{}_{Q^{\prime}}}\operatorname{sech}\rho_{b}+z_{c_{2}}}\right]\right),$ (46)
where the conformal factors $\Omega_{P^{\prime}}$ and $\Omega_{Q^{\prime}}$ are now given by eq.˜43, and have once again assumed $P^{\prime}\equiv(-y_{{}_{P^{\prime}}},\tau_{{}_{P^{\prime}}})$ and $Q^{\prime}\equiv(-y_{{}_{Q^{\prime}}},\tau_{{}_{Q^{\prime}}})$. Due to the presence of the island region on the brane, the contributions due to the area term in eq.˜8 must also be taken into consideration. This term may be directly obtained from Deng:2023pjs as
$S_{area}$ $=\frac{1}{4G_{N}}\cosh^{-1}\left[\frac{\sqrt{2z_{c_{1}}y_{{}_{P^{\prime}}}\operatorname{sech}\rho_{b}+y_{{}_{P^{\prime}}}^{2}+z_{c_{1}}^{2}}}{y_{{}_{P^{\prime}}}\operatorname{sech}\rho_{b}+z_{c_{1}}}\right]$
$\qquad\qquad+\frac{1}{4G_{N}}\cosh^{-1}\left[\frac{\sqrt{2z_{c_{2}}y_{{}_{Q^{\prime}}}\operatorname{sech}\rho_{b}+y_{{}_{Q^{\prime}}}^{2}+z_{c_{2}}^{2}}}{y_{{}_{Q^{\prime}}}\operatorname{sech}\rho_{b}+z_{c_{2}}}\right].$ (47)
@@
$S(A)$ $=\frac{c}{6}\left(\log\frac{2x_{1}}{z_{c_{1}}}+\log\frac{2x_{2}}{z_{c_{2}}}+2\rho_{b}+2\log\frac{2\ell}{\epsilon_{y}\operatorname{sech}\rho_{b}}\right)-\frac{c}{6}\frac{z_{c_{1}}e^{\rho_{b}}}{x_{1}}-\frac{c}{6}\frac{z_{c_{2}}e^{\rho_{b}}}{x_{2}}$
$=\frac{c}{3}\left(\rho_{c}+\rho_{b}+\log\frac{2\ell}{\epsilon_{y}\operatorname{sech}\rho_{b}}\right)-\frac{2c}{3}e^{\rho_{b}-\rho_{c}},$ (48)
where in the final step we utilize the transformations in eqs.˜21, 23 and 25 to express the results in the AdS black string coordinates. Note that in the limit $\rho_{c}\to\rho_{\epsilon}$, the dominant term (the first term in section˜4.2.1) is in exact agreement with the results of the undeformed scenario given by section˜3.2.1.
Bulk computation : From the bulk perspective the entanglement entropy may be determined using the DES formula in eq.˜9, where the first term may be evaluated in terms of the geodesic lengths of the two RT surfaces $PP^{\prime}$ and $QQ^{\prime}$ using eq.˜32 as
$S_{RT}=$ $\frac{1}{4G_{N}}\cosh^{-1}\left[\frac{(x_{1}+y_{{}_{P^{\prime}}}\tanh\rho_{b})^{2}+(z_{c_{1}}+y_{{}_{P^{\prime}}}\operatorname{sech}\rho_{b})^{2}+(\tau_{1}-\tau_{{}_{P^{\prime}}})^{2}+z_{c_{1}}^{2}}{2z_{c_{1}}(z_{c_{1}}+y_{{}_{P^{\prime}}}\operatorname{sech}\rho_{b})}\right]$
$\qquad+\frac{1}{4G_{N}}\cosh^{-1}\left[\frac{(x_{2}+y_{{}_{Q^{\prime}}}\tanh\rho_{b})^{2}+(z_{c_{2}}+y_{{}_{Q^{\prime}}}\operatorname{sech}\rho_{b})^{2}+(\tau_{2}-\tau_{{}_{Q^{\prime}}})^{2}+z_{c_{2}}^{2}}{2z_{c_{2}}(z_{c_{2}}+y_{{}_{Q^{\prime}}}\operatorname{sech}\rho_{b})}\right],$ (49)
where we now utilize eq.˜41 to express the coordinates of the brane points as $P^{\prime}\equiv(\tau_{{}_{P^{\prime}}},-y_{{}_{P^{\prime}}}\tanh\rho_{b},z_{c_{1}}+y_{{}_{P^{\prime}}}\operatorname{sech}\rho_{b})$ and $Q^{\prime}\equiv(\tau_{{}_{Q^{\prime}}},-y_{{}_{Q^{\prime}}}\tanh\rho_{b},z_{c_{2}}+y_{{}_{Q^{\prime}}}\operatorname{sech}\rho_{b})$. The defect term in eq.˜9 contributes in this phase due to the presence of entanglement entropy islands on the brane, which may be directly obtained from Deng:2023pjs . The generalized entanglement entropy may then be determined as
$S_{gen}=S_{eff}+\frac{c}{6}\log\frac{2y_{{}_{P^{\prime}}}}{z_{c_{1}}(z_{c_{1}}+y_{{}_{P^{\prime}}}\operatorname{sech}\rho_{b})}+\frac{c}{6}\log\frac{2y_{{}_{Q^{\prime}}}}{z_{c_{2}}(z_{c_{2}}+y_{{}_{Q^{\prime}}}\operatorname{sech}\rho_{b})}.$ (50)
Extremizing the above expression we obtain, upto first order in $z_{c_{1}}$ and $z_{c_{2}}$, $\tau_{{}_{P^{\prime}}}=\tau_{1},y_{{}_{P^{\prime}}}=x_{1}-2z_{c_{1}}e^{\rho_{b}}$ and $\tau_{{}_{Q^{\prime}}}=\tau_{2},y_{{}_{Q^{\prime}}}=x_{2}-2z_{c_{2}}e^{\rho_{b}}$. The final holographic entanglement entropy may then be obtained as section˜4.2.1 on application of the transformations in eqs.˜21, 23 and 25 and the Brown-Henneaux relation.
@@ 4.2.2 No-Island Phase
Field theory computation : This phase corresponds to the scenario where the points $P$ and $Q$ are located far away from the EOW brane, as a result of which the radiation subsystem $A$ has no corresponding entanglement entropy island on the brane, as illustrated in fig.˜5. Consequently, the area term in eq.˜8 does not contribute and the entanglement entropy is equivalent to the effective term in this case, which may be determined in terms of the 2-point correlator of twist fields at $P$ and $Q$ as given in eq.˜36. The rest of the computation follows as elaborated in section˜3.2.2, and the final entanglement entropy may then be expressed as
$S(A)=\frac{c}{6}\log\left[\frac{u_{h}}{LR}\left(\Delta_{L}+\Delta_{R}+2\sqrt{\Delta_{L}\Delta_{R}}\cosh\frac{t}{u_{h}}\right)\right]+\frac{c}{3}\rho_{c}.$ (51)
@@
Bulk computation : Utilizing the DES formula in eq.˜9, the holographic entanglement entropy in the no-island phase is simply given by the first term. As elaborated earlier, due to the absence of entanglement entropy islands on the EOW brane the defect term does not contribute in this scenario. The holographic entanglement entropy, which in this case is simply $S_{RT}$, is thus obtained as
$S(A)=S_{RT}=\frac{1}{4G}\cosh^{-1}\left[\frac{(x_{2}-x_{1})^{2}+(\tau_{2}-\tau_{1})^{2}+z_{c_{1}}^{2}+z_{c_{2}}^{2}}{2z_{c_{1}}z_{c_{2}}}\right],$ (52)
which matches with eq.˜51 on application of the transformations described in eqs.˜21, 23 and 41 and the Brown-Henneaux relation.
@@ 5 Page curve
In this section we plot and analyse the Page curves for the entanglement entropy of the radiation subsystem $A$, for both undeformed and $T\bar{T}$ deformed scenarios. To this end, we evaluate the entanglement entropy under the assumption $c=6,u_{h}=1,\epsilon_{y}=0.01,L=R=0.5$, and $\rho_{\epsilon}=\rho_{c}=4$. [^6] We assume that $\rho_{c}$ and $\rho_{\epsilon}$ to have same values in this scenario. This is because the objective of this analysis is to investigate the changes in the original Page curve when corrections due to $T\bar{T}$ deformation is taken into account. For this purpose we require the undeformed entanglement entropy to have the same value numerically for both scenarios. In addition we systematically vary $\rho_{b}$ to observe how the Page curve changes with the brane angle. This is illustrated in fig.˜6, where the Page curves corresponding to distinct brane angles are depicted using different colors. The Page curve for the deformed scenarios are depicted by the solid curves, while those for the undeformed scenarios are depicted by the dashed curves. We observe that for both scenarios the entanglement entropy increases with time in the no-island phase. Upon transition to the island phase the entanglement entropy stabilizes to a constant value, which varies with change in the brane angle.
A general trend emerges from fig.˜6. We observe in the no-island phase the Page curves overlap for both undeformed and $T\bar{T}$ deformed scenarios and for all values of $\rho_{b}$, indicating that the evolution of the entanglement entropy is independent of the deformation parameter or the brane angle in this phase. However, we observe some interesting features in the island phase. The entanglement entropy in this phase is dependent on both the brane angle and the deformation parameter, with the deformed entanglement entropy being smaller than the undeformed one for a given brane angle. Moreover, the Page time $T_{p}$ (the time of transition between the no-island and the island phases of the entanglement entropy) appears to be dependent on the brane angle as well as the deformation parameter. This dependence can be mathematically explored by deriving the expressions of $T_{p}$ for both the undeformed and the $T\bar{T}$ deformed scenarios as
$T_{p}=\begin{cases}u_{h}\cosh^{-1}\left[\frac{LR\left(e^{2\rho_{b}}+1\right)^{2}-u_{h}\epsilon_{y}^{2}(\Delta_{L}+\Delta_{R})}{2u_{h}\epsilon_{y}^{2}\sqrt{\Delta_{L}\Delta_{R}}}\right]\qquad&\text{Undeformed scenario,}\\ u_{h}\cosh^{-1}\left[\frac{LR\left(e^{2\rho_{b}}+1\right)^{2}e^{-4e^{\rho_{b}-\rho_{c}}}-u_{h}\epsilon_{y}^{2}(\Delta_{L}+\Delta_{R})}{2u_{h}\epsilon_{y}^{2}\sqrt{\Delta_{L}\Delta_{R}}}\right]\qquad&T\bar{T}\text{-deformed scenario,}\end{cases}$ (53)
@@
It is interesting to further investigate the dependence of the Page time $T_{p}$ on the brane angle and the deformation parameter and explore any new physics that might emerge. We explore the dependence of $T_{p}$ on the brane angle in fig.˜7 for both the undeformed and the $T\bar{T}$ deformed scenarios. Note that though the Page time initially increases with $\rho_{b}$, it rapidly drops to zero for the $T\bar{T}$ deformed scenarios as $\rho_{b}$ further increases. This is in contrast with the undeformed scenario, where the Page time increases consistently with $\rho_{b}$. This behaviour is also observed in Basu:2024xjq , where the authors speculate that this rapid decrease in the Page time might be due the CFT being no longer located at the asymptotic boundary at $\rho\to\infty$, but pushed inwards to a finite value in the presence of $T\bar{T}$ deformation.
@@
We further extend the analysis by examining how the Page time changes with the deformation parameter in fig.˜8, where we plot $T_{p}$ against $e^{-\rho_{c}}$ (related to the deformation parameter as given in eq.˜45) for different brane angles. $e^{-\rho_{c}}\to 0$ corresponds to the undeformed scenario, while increasing values indicate $T\bar{T}$ deformed scenarios. We observe that for all brane angles the Page time decreases approximately linearly with the UV cut-off $e^{-\rho_{c}}$, but shows a sudden drop at the end. With increase in $\rho_{b}$, $T_{p}$ starts at a higher value, but decreases much rapidly for high brane angles compared to low brane angles.
@@ 6 Summary and Discussions
In this article, we investigate the proposed duality between the island and the defect extremal surface (DES) prescription for the fine-grained entanglement entropy in a Karch-Randall (KR) braneworld model, where the radiation baths are located in a gravitational background. In particular, we consider an AdS3 black string geometry truncated by an end-of-the-world (EOW) brane for which the effective lower-dimensional perspective consists of a radiation bath defined on an AdS2 black hole background. We consider subsystems on radiation baths corresponding to AdS2 eternal black holes in the lower-dimensional effective picture, and compute the entanglement entropy for distinct phases using the island formula in the effective $2d$ and the DES formula in the bulk picture. We demonstrate an agreement between the two prescriptions, thus establishing that the DES formula is a holographic dual to the island formula. Additionally, the results agree with those existing in the literature obtained using the bulk and the boundary prescription Geng:2021mic ; Geng:2022dua , which acts as a check for the accuracy of our computations.
We further investigate the entanglement structure for AdS3 black string geometry with a finite cut-off arising from $T\bar{T}$ deformation McGough:2016lol , truncated by an EOW brane. The usual AdS/CFT duality applies to the Dirichlet boundary on the cut-off surface, while Neumann boundary conditions apply on the EOW brane Deng:2023pjs . We consider an eternal AdS2 black hole coupled to radiation baths (now located on a finite cut-off surface induced by $T\bar{T}$ deformation), and determine the entanglement entropy up to first order for radiation subsystems using both the island and the DES formulas. We observe an agreement between the two prescriptions for both the island and no-island phases of the entanglement entropy, providing another strong consistency check for the proposed duality.
@@
Though Chen:2018eqk ; Murdia:2019fax ; He:2019vzf have established that modifying the correlator by appropriate insertions of stress tensor components correctly captures the effects of the $T\bar{T}$ deformation, in this article we adopt a different method for our analysis. Similar to Deng:2023pjs , we assume that this deformation manifests through modifications in the induced metric defined on the radial cut-off surface and observe an agreement between the results obtained using the island and the DES prescriptions. However, a perturbative theory of $T\bar{T}$ deformed BCFT must still be developed.
This work identifies several promising directions to explore further. An important consideration is analysing the mixed state entanglement structure for both undeformed and deformed scenarios for KR braneworld models. Extending the framework to higher dimensions is an equally crucial advancement, particularly in light of Chu:2021gdb , which shows that the DES and the island prescriptions do not align consistently for higher dimensions even for undeformed scenarios. Furthermore, testing the proposed duality between these prescriptions for other deformed theories is also a compelling direction. We leave these open questions for future study.